using the sequential definition of continuity, prove that the function f :r to r , defined by f (x) = 3x^2+7, for all x belong to r, is continuous
1
Expert's answer
2020-05-18T10:18:43-0400

Sequential definition of continuity states if f is continuous at a if and only if "f(x_n) \to f(a)" for all sequences "x_n \to a."

Given function is "f (x) = 3x^2+7".

Let "{<x_n>}" be any sequence convergence to "a".

Now, "f(x_n) - f(a) = (3x_n^2+7) - (3a^2+7) = 3(x_n^2-a^2)"

So, "|f(x_n)-f(a)|= |3(x_n+a)(x_n-a)| \leq 3|x_n+a||x_n-a|".

Now, we have "x_n\to a" so "{x_n}" is bounded sequence "\implies |x_n|\leq k \hspace{0.05 in} \forall \hspace{0.05 in} n \geq m" for some finite m.

"\implies |x_n+a|\leq k+a" for all "n\geq m." .

So, "|f(x_n)-f(a)| \leq 3|x_n+a||x_n-a| \leq 3(k+a) |x_n-a|" for all "n\geq m" .

Thus, as "x_n \to a", "f(x_n) \to f(a)."


So, by Sequential definition of continuity, f is continuous at every real number.